JÓZSEF ATTILA UNIVERSITY  Aradi vértanúk tere 1.

Szeged, Hungary–6720

BOLYAI INSTITUTE  Tel/Fax: (36) 62 326246

Conference on Lattices and Universal Algebra

Szeged, Hungary (August 3–7, 1998)

SECOND (LAST) ANNOUNCEMENT

The conference will start at 10 a.m. on August 3 (Monday), and will conclude on August 7 (Friday) early afternoon. The participants are expected to arrive on August 2. The lectures will take place at the Bolyai Institute. The scientific program will consist of 50minute invited plenary lectures by Brian Davey, George Grätzer, Victor Gorbunov, Christian Herrmann, Emil W. Kiss, Ralph McKenzie, Matthew Valeriote, Friedrich Wehrung and Ross Willard, and numerous 20-minute talks in parallel sections. Those wishing to use an overhead projector for their talks are requested to bring ready-made transparencies. The list of participants, the title (and soon the schedule) of their talks and some other pieces of information are available at our web site http://www.math.u-szeged.hu/ (click on Conferences). This web site will be updated from time to time, so please inform us on changes in the status of your talk (like on the new value of ”to be announced later”). The banquet of the conference is scheduled for Monday evening. The problem session will be combined with a small excursion to Hódmezövásárhely^{†} on Wednesday. Although the deadline for submitting papers for the proceedings, a separate issue of Algebra Universalis, is September 30, we encourage to submit papers during the conference.

Travel information

How to travel in Budapest. All trams, buses, trolley buses and the metro (underground) need exactly the same pre-purchased tickets, which you have to validate yourself. These tickets are sold (even on Sundays) at newspaper stands at railway stations, metro stations and at both airports on the departure level. The price of a ticket is 75 Ft (seventy five Forints). The current rate is about 1 US dollar = 215 Ft, this may slightly vary. Exchange offices are open every day at the main railway stations and at the airports. The three main railway stations are called Nyugati (pu.), Keleti (pu.) and Déli (pu.). These and Köbánya–Kispest are connected by metro lines. In details, the blue metro line (indicated by blue letter M or M3) connects Nyugati and Köbánya–Kispest, the red metro line (M2) connects Keleti and Déli, and these two lines meet at Deák tér (where you have to validate another ticket when changing M2 to M3 or conversely). Thirty minutes should suffice to reach any railway station from any other one.

There is a bus (line 93, red number!!!) operating between the airports and Köbánya– Kispest. It runs every 20 minutes and needs the above-mentioned ticket. At the airports (called Ferihegy 1 and Ferihegy 2) you have to choose the direction ”center”. Then, after 25 minutes or so, Köbánya–Kispest is the last stop, and going upstairs (where you see a lot of people coming and going) you can easily find the metro (and also the railway) station.

† To avoid TEX problems at some computers, sometimes we just approximate the correct Hungarian spelling of accents.

As a comfortable possibility to reach the railway station Nyugati (or any other spot in Budapest) from the airports, you can make use of the shuttle-bus (airport mini bus) advertised at the baggage claim. It costs approximately 1300 Ft. We do not recommend to take a taxi; some drivers may overcharge passengers (especially the foreign ones).

Trains. The train connection between Budapest and Szeged is good. Below please find the time tables, the first one from Budapest to Szeged, and the other one from Szeged to Budapest. At departure and arrival times we use the abbreviations Bp for Budapest Nyugati and Sz for Szeged. The meaning of other notations is given later. Each pair of brackets represents a train. All trains for Szeged depart from Nyugati, and the non-IC trains stop at Köbánya–Kispest (about a 13 minute travel from Nyugati).

\underline{Trains from Budapest Nyugati to Szeged}: [Bp 6.25 − Sz 9.01], [Bp 7.15 − Sz 9.27 (IC Napfény)], [Bp 8.15 − Sz 10.59 (+Miskolc!)], [Bp 10.25 − Sz 13.08], [Bp 12.25 − Sz 14.59], [Bp 14.25 − Sz 17.12], [Bp 15.15 − Sz 17.27 (IC Dankó Pista)], [Bp 16.25 − Sz 18.59], [Bp 18.20 − Sz 21.02], [Bp 19.15 − Sz 21.27 (IC Maros)].

\underline{Trains from Szeged to Budapest Nyugati}: [Sz 5.20 − Bp 7.55 (+Miskolc!)], [Sz 6.30 − Bp 8.43 (IC Napfény)], [Sz 6.50 − Bp 9.32], [Sz 9.35 − Bp 12.10], [Sz 11.15 − Bp 13.53], [Sz 13.20 − Bp 15.54], [Sz 14.30 − Bp 16.43 (IC Dankó Pista)], [Sz 15.20 − Bp 17.58 (+Miskolc!)], [Sz 17.20 − Bp 19.54], [Sz 18.30 − Bp 20.43 (IC Maros)].

The one way price (at present) is 1710 Ft (1st class) or 1140 Ft (2nd class). For IC trains a seat reservation is obligatory (290 Ft). The names of IC trains are also given. In the direction Budapest — Szeged, Szeged is the last stop. When ”+Miskolc!” is indicated then be careful: some carriages (usually the last ones of the train) go to Miskolc instead of Szeged or Budapest.

Arriving by car. You can reach Szeged via the National Highway no. 5 (Budapest– Szeged), which is identical to a section of the European Highway E75. Although the highway prices (about 1600 Ft for a single trip) look normal, they are the highest in Europe if we count them per kilometer. If you enter Hungary from Austria at Hegyeshalom, then you pay the price only for less than fifty kilometers, not up to Budapest. When you drive between Budapest and Szeged, then you have to pay again but only for the first half (Budapest–Kiskunfélegyháza). In both cases it is possible to choose a free (but slower) way. Parking in the downtown area in Szeged on week days is possible only with prepurchased parking tickets. The Bolyai Institute is located at the boarder of the downtown area.

Getting around Szeged. Trams, buses and trolley buses operate with the same tickets, which are different from those valid in Budapest. A ticket costs 65 Ft if bought in advance (at the railway station or, e.g., in Aradi vértanúk tere), but you can also buy it at the price 90 Ft after boarding the vehicle. Daily and weakly tickets are available at the price 240 Ft and 710 Ft, respectively. The conference will take place at the Bolyai Institute (mathematics building of the university), on the eastern (right hand) side of Aradi vértanúk square. You can reach this square by taking the (only) tram in front of the railway station (in the only direction), then Aradi vértanúk square is the third stop. Equivalently, it is the first square or it is the first stop after the tram goes through the Gate of Heroes. Walking along the tram lines needs twelve minutes or so. Sunday afternoon some of our colleagues will try to meet and assist you at Szeged Railway Station. They are recognizable by an issue of Algebra Universalis in their hands.

Registration. Upon arrival to Szeged you are kindly requested to come in the registration office at Bolyai Institute, Aradi vértanúk square. On August 2 (Sunday) the registration office will operate between 3 p.m. to 10 p.m. Some additional registration hours will be provided in the next two days.

Mini dictionary. We list some words and two typical sentences. If — being uncertain about pronunciation — you write them on a piece of paper, do not omit the accents.

please (give me)  =  kérek
ticket, metro ticket, bus ticket  =  jegy, metrojegy, buszjegy
train ticket, return ticket  =  vonatjegy, retúr jegy
1st class, 2nd class  =  1. osztályú, 2. osztályú
to Szeged, to Budapest  =  Szegedre, Budapestre
IC seat reservation [departure time/name of train]  =  IC helyjegy [...]
9.25 a.m., 2.30 p.m.  =  9.25, 14.30
nonsmoking, smoking  =  nemdohányzó, dohányzó
for August 6, 7, 8, 9  =  Augusztus 6-ra, 7-re, 8-ra, 9-re

A first class ticket to Szeged, please = Egy 1. osztályú jegyet kérek Szegedre. (And) please give me a nonsmoking IC seat reservation to Dankó Pista that departs at 15.15 = (És) kérek egy nemdohányzó helyjegyet a 15.15-kor induló Dankó Pistára.

Szeged, July 2, 1998.

G. Czédli  E. K. Horváth  G. Takách  L. Zádori

(the organizing committee)